OpenStudy (anonymous):

What is the net force needed to accelerate a full grocery sack weighing 210 N upward at a rate of 1.5 m/s2?

OpenStudy (loser66):

 F - W = m a find m, m = W/g = 210/9.8 = 21.43kg F= ma +W F = 21.43(kg)* 1.5 (m/s^2) + 210 N = 242.145N = 2.4 *10^2 N (2 significant figure)